### v1

@@ Appendix F Additional Analysis
F.1 Number of steps taken in each stage by different baselines
@@
Heuristic Heuristic Ground Truth 0485.3 000- 0095.9 0008.5 0574.1
Heuristic RL Ground Truth 0483.5 0007.7 0101.9 0067.2 0611.6
RL Heuristic Ground Truth 0313.4 000- 0136.9 0007.7 0437.6
RL RL Ground Truth 0327.6 0009.1 0130.0 0047.8 0433.6
Heuristic Heuristic DETIC [26] 0405.9 000- 0048.9 0006.8 0459.0
Heuristic RL DETIC [26] 0412.8 0044.7 0047.5 0242.2 0584.2
RL Heuristic DETIC [26] 0504.3 000- 0128.4 0007.3 0586.2
RL RL DETIC [26] 0496.3 0045.9 0139.0 0156.4 0583.3
@@
Heuristic Heuristic Ground Truth 100.0 00- 055.2 055.2 048.9 048.8
Heuristic RL Ground Truth 100.0 054.7 053.7 053.7 046.7 036.3
RL Heuristic Ground Truth 100.0 00- 080.6 080.6 071.2 071.1
RL RL Ground Truth 100.0 079.5 068.1 068.1 060.2 048.2
Heuristic Heuristic DETIC [26] 100.0 00- 031.8 031.8 027.8 027.8
Heuristic RL DETIC [26] 100.0 032.3 017.6 017.6 015.5 004.3
RL Heuristic DETIC [26] 100.0 00- 050.2 050.2 037.3 037.2
RL RL DETIC [26] 100.0 050.1 024.8 024.8 019.8 008.5
@@
F.2 Performance on Seen vs. Unseen Object Categories
@@
Heuristic Heuristic Ground Truth 047.9 042.3 046.2 040.3 038.0 039.5 018.2 019.7 018.6 006.3 004.9 005.9
Heuristic RL-PPO Ground Truth 047.9 045.3 047.2 042.0 041.0 041.7 026.5 028.6 027.1 010.7 011.9 011.0
RL-PPO Heuristic Ground Truth 054.6 056.1 055.1 042.5 040.7 041.9 025.9 027.2 026.4 004.9 005.9 005.2
RL-PPO RL-PPO Ground Truth 055.5 055.8 055.7 050.5 049.3 050.2 035.6 034.2 035.2 012.3 0010.0 011.6
Heuristic Heuristic DETIC [26] 023.6 022.6 023.3 011.7 011.1 011.5 002.9 003.2 003.0 000.2 000.5 000.3
Heuristic RL-PPO DETIC [26] 025.5 022.9 024.8 0010.0 008.4 009.5 005.4 004.3 005.0 000.2 000.3 000.2
RL-PPO Heuristic DETIC [26] 020.2 019.1 019.9 010.5 0010.0 010.2 004.7 003.8 004.4 000.9 000.0 000.6
RL-PPO RL-PPO DETIC [26] 020.0 019.1 019.8 012.2 011.1 011.8 007.1 004.9 006.3 000.9 000.8 000.8
@@
F.2.1 Example DETIC [26] predictions
In Table 5, we observe that Gaze policy takes significantly longer time to terminate with DETIC [26] perception. The gaze policy (see Fig. 17) tries to center the agent on the object of interest by allowing the agent to move its base and camera tilt. For this, it relies on DETIC’s ability to detect novel objects. Now, we visualize DETIC segmentations of agent’s egocentric observations by placing agent at the points where the Gaze skill is expected to start: the object’s viewpoints. We observe that while DETIC succeeds in a few cases, it fails at consistently detecting the objects in the egocentric frame.
@@ F.2.1 Example DETIC [26] predictions
Appendix G Hardware Setup
@@ G.1 Hardware Choice
We describe some options for commercially-available robotics hardware in Tab. 8. While the Franka Emika Panda is not a mobile robot, we include it here because it’s a very commmonly used platform in both industrial research labs and at universities, making its price a fair comparison point for what is reasonable.
G.2 Robot Setup
@@
Figure 19: Network setup diagram for HomeRobot. We can run visualizations on a GPU-enabled workstation, while running only the necessary code on a robot for low-level control and SLAM.
@@
Timing between the robot and the remote workstation. We use ROS [105] as our communications layer, and to implement low-level control on the robot. This also provides network communication. However, due to potential latency between the robot and desktop, we also need to make sure that observations are up to date.
We set up the robot to block after executing most navigation motions, in order to make this process simpler, until there is an up to date image observation from the robot side. This means that timing between the robot and the workstation is extremely important: if we do not have up to date timing, we might have SLAM poses and depth measurements that do not match, which will lead to worse performance.
We solved this by having a clock on the robot side publish its time over ROS, and configure all systems to use this ROS master clock instead of system time. This prevents the user from having to worry about Linux time synchronization protocols like NTP when setting up the robot for the first time.
G.3 Visualizing The Robot
@@
We use RVIZ, a part of ROS, to visualize results and progress. Fig. 20 shows three different outputs from our system: on the far left, an image from the test environment being processed by Detic; in the center, a top-down map generated by the navigation planner described in Sec. D.2; and on the right, an image from RVIZ with the point cloud from the robot’s head camera registered against the 2D lidar map created by Hector SLAM.
One advantage of the HomeRobot stack is that it is designed to work with existing debugging tools - especially ROS [105]. ROS is a widely-used framework for robotics software development which comes with a lot of online resources, official support from Hello Robot, and a rich and thriving open-source community with wide industry backing.
G.4 Using The Stretch: Navigation vs. Position Mode
@@
  • Navigation mode: the robot’s camera is facing forward; we use reactive low-level control for navigation; robot can rotate in place, roll backwards, and will reactively track goals sent from the desktop.

### v2

@@ Appendix G Additional Analysis
G.1 Number of steps taken in each stage by different baselines
@@
Heuristic Heuristic Ground Truth 0291.8 000- 0065.5 0008.4 0360.5
Heuristic RL Ground Truth 0293.5 0019.4 0064.3 0084.4 0438.7
RL Heuristic Ground Truth 0295.1 000- 0105.0 0007.0 0401.6
RL RL Ground Truth 0302.4 0025.7 0112.8 0045.9 0455.2
Heuristic Heuristic DETIC [27] 0335.0 000- 0029.5 0006.7 0361.8
Heuristic RL DETIC [27] 0330.1 0152.0 0027.5 0068.2 0556.5
RL Heuristic DETIC [27] 0509.5 000- 0153.3 0007.1 0610.4
RL RL DETIC [27] 0539.1 0101.3 0124.4 0033.7 0634.7
@@
Heuristic Heuristic Ground Truth 100.0 000- 065.1 065.1 062.1 062.1
Heuristic RL Ground Truth 100.0 065.6 064.3 064.3 061.3 052.2
RL Heuristic Ground Truth 100.0 000- 076.3 076.2 066.8 066.8
RL RL Ground Truth 100.0 077.0 074.5 074.5 065.1 060.6
Heuristic Heuristic DETIC [27] 100.0 000- 034.7 034.7 031.1 031.1
Heuristic RL DETIC [27] 100.0 033.9 027.2 027.2 024.4 017.6
RL Heuristic DETIC [27] 100.0 000- 032.9 032.7 024.2 024.2
RL RL DETIC [27] 100.0 034.7 024.9 024.8 018.1 015.3
@@
G.2 Performance on Seen vs. Unseen Object Categories
@@
Heuristic Heuristic Ground Truth 050.9 053.2 054.1 046.4 047.2 048.5 027.5 030.0 031.5 004.1 005.2 005.1
Heuristic RL Ground Truth 054.9 058.2 056.5 048.6 055.1 051.5 039.0 037.3 042.3 014.5 012.0 013.2
RL Heuristic Ground Truth 067.1 064.1 065.4 055.0 051.0 054.8 044.8 044.8 043.7 006.2 007.8 007.3
RL RL Ground Truth 068.4 065.8 066.6 063.7 057.6 061.1 054.7 049.0 050.9 015.7 014.4 014.8
Heuristic Heuristic DETIC [27] 022.2 021.1 028.7 012.5 010.5 015.2 003.2 003.3 005.3 000.9 000.7 000.4
Heuristic RL DETIC [27] 022.4 022.2 029.4 011.9 011.8 013.2 005.1 003.5 005.8 000.3 001.4 000.5
RL Heuristic DETIC [27] 018.7 023.0 021.9 009.9 011.8 011.5 005.8 005.3 006.0 000.3 000.0 000.6
RL RL DETIC [27] 021.5 020.7 021.7 010.9 011.0 010.2 006.9 006.2 006.2 001.0 000.7 000.4
@@
G.2.1 Example DETIC predictions
In Table 5, we observe that the Gaze policy takes a significantly longer time to terminate with DETIC [27] perception. The gaze policy (see Fig. 19) tries to center the agent on the object of interest by allowing the agent to move its base and camera tilt. For this, it relies on DETIC’s ability to detect novel objects. Now, we visualize DETIC segmentations of agent’s egocentric observations by placing agent at the points where the Gaze skill is expected to start: the object’s viewpoints. We observe that while DETIC succeeds in a few cases, it fails at consistently detecting the objects in the egocentric frame.
@@ G.2.1 Example DETIC predictions
Appendix H Hardware Setup
@@ H.1 Hardware Choice
We describe some options for commercially available robotics hardware in Tab. 8. While the Franka Emika Panda is not a mobile robot, we include it here because it’s a very commonly used platform in both industrial research labs and at universities, making its price a fair comparison point for what is reasonable.
H.2 Robot Setup
@@
Figure 21: Network setup diagram for HomeRobot. We can run visualizations on a GPU-enabled workstation while running only the necessary code on a robot for low-level control and SLAM.
@@
Timing between the robot and the remote workstation. We use ROS [113] as our communication layer, and to implement the low-level control on the robot. This also provides network communication. However, due to potential latency between the robot and the desktop, we also need to make sure that observations are up to date.
We set up the robot such that it stops after executing most navigation actions, until there is an up to date image observation from the robot side. This means that time synchronization between the robot and the workstation is extremely important; if we do not have up-to-date timing, we might have SLAM poses and depth measurements that do not match, which will lead to worse performance.
We solved this by having a clock on the robot side publish its time over ROS, and configuring all systems to use this ROS master clock instead of system time. This prevents the user from having to worry about Linux time synchronization protocols like NTP when setting up the robot.
H.3 Visualizing The Robot
@@
We use RVIZ, a part of ROS, to visualize results and progress. Fig. 22 shows three different outputs from our system: on the far left, an image from the test environment being processed by Detic; in the center, a top-down map generated by the navigation planner described in Sec. E.2; and on the right, an image from RVIZ with the point cloud from the robot’s head camera registered against the 2D lidar map created by Hector SLAM.
One advantage of the HomeRobot stack is that it is designed to work with existing debugging tools - especially ROS [113]. ROS is a widely-used framework for robotics software development that comes with a lot of online resources, official support from Hello Robot, and a rich and thriving open-source community with wide industry backing.
H.4 Using The Stretch: Navigation vs. Position Mode
@@
  • Navigation mode: the robot’s camera is facing forward; we use reactive low-level control for navigation; the robot can rotate in place, roll backward, and will reactively track goals sent from the desktop.